# Adversarial Attacks to Latent Representations of Distributed Neural Networks in Split Computing

### v4

@@ Abstract
Distributed deep neural networks (DNNs) have been shown to reduce the computational burden of mobile devices and decrease the end-to-end inference latency in edge computing scenarios. While distributed DNNs have been studied, to the best of our knowledge, the resilience of distributed DNNs to adversarial action remains an open problem. In this paper, we fill the existing research gap by rigorously analyzing the robustness of distributed DNNs against adversarial action. We cast this problem in the context of information theory and rigorously proved that (i) the compressed latent dimension improves the robustness but also affect task-oriented performance; and (ii) the deeper splitting point enhances the robustness but also increases the computational burden. These two trade-offs provide a novel perspective to design robust distributed DNN. To test our theoretical findings, we perform extensive experimental analysis by considering 6 different DNN architectures, 6 different approaches for distributed DNN and 10 different adversarial attacks using the ImageNet-1K dataset.
@@
Distributed DNN , Adversarial Robustness , Split Computing
@@ 1 Introduction
Deep neural networks (DNNs) have achieved significant success in various domains such as computer vision [1], natural language processing [2], and wireless communication [3], among many others. However, state-of-the-art DNNs are challenging to deploy on resource-limited mobile devices. While mobile-specific DNNs have been proposed [4], they usually come with a significant loss in accuracy. On the other hand, completely offloading the computation to edge or cloud computers is impractical in mobile scenarios due to the excessive communication overhead corresponding to the transfer of the DNN input from the mobile device to the edge/cloud [5]. A new paradigm called distributed computing – also referred to as split computing in prior art [6] – divides the computation of DNNs across multiple devices – according to the available processing power and networking bandwidth. The key advantage is that optimal load distribution can be achieved while meeting maximum end-to-end latency constraints and also preserving the DNN accuracy [7]. For an excellent survey on distributed/split computing, the reader is referred to the work by [6].
Without loss of generality, we assume that a DNN model is divided into a mobile DNN and a local DNN, respectively executed by the mobile device and an edge/cloud computer. Usually, the DNN architecture is modified by introducing a compression layer ( “bottleneck”) at the end of the mobile DNN [8, 9, 10, 11, 12], which is trained to learn a latent representation that reduces the amount of data being sent to the edge/cloud. The compressed representation is then used by the local DNN to produce the final prediction output (e.g., classification).
@@
Despite its significance and timeliness, to the best of our knowledge, assessing the robustness of distributed DNNs remains an unexplored problem. We remark that achieving a fundamental understanding of these attacks and evaluating their effectiveness in state-of-the-art DNNs is paramount to design robust distributed DNNs. To this end, we theoretically analyze the robustness of distributed DNNs using Information Bottleneck (IB) theory [13]. Our first key theoretical finding is that the latent robustness is highly depended on the depth of splitting point. With similar levels of information distortion, latent representations in deeper layers exhibit constantly better robustness compare to those in early layers. Our second key finding is that the trade-off between DNN performance and robustness is intrinsically related to the dimension of the latent space. While a smaller latent space may increase robustness by reducing the model variance, it will also introduce bias in the model thus affecting the generalization capability of the DNN model.
We extensively evaluate our theoretical findings by considering 10 adversarial algorithms, i.e., 4 white-box attacks [14, 15, 16, 17] and 6 black-box attacks [18, 19, 20, 21, 22, 23]. We apply these attacks to 6 reference architectures [24, 25] designed with 6 distributed DNN approaches [26, 11, 12, 7, 27, 28]. The experimental results validate our theoretical findings on the examined DNNs and attack algorithms.
@@
$\bullet$ We perform extensive experiments with the ImageNet-1K [30] dataset, by considering 6 different DNN architectures, 6 different distributed DNN approaches under 10 different attacks to support our theoretical findings. The results show that the theoretical analysis applies to the experimental settings under consideration. We share our code for reproducibility at https://github.com/asdfqwezxcf/AdvLatent, and we hope this work may open the door to a new field dedicated to studying the resilience of distributed DNNs.
@@ 2 Threat Model
where $||\delta_{d}||_{p}\leq\sigma$ and $\sigma$ is the distance constraint under different $l_{p}$ norm. Additionally, for visual applications, $\delta_{d}$ should satisfy the condition $x+\delta_{d}\in[0,1]^{d}$ as there is an explicit upper and lower bound for red, green and blue (RGB) value in digital images.
@@
where $||\delta_{t}||_{p}\leq\sigma$ is the distance constraint under $l_{p}$ norm. We remark that the latent representations are model-dependent and there is no explicit bound for their value other than their computer-level representation (e.g., float, integer, double).
@@ 3.1 Analysis through Information Bottleneck
The end-to-end robustness of the DNN model can be modeled with $I(Y;\hat{Y})$, which measures the mutual information between the label $Y$ and the DNN inference $\hat{Y}$ [34]. We apply the Data Processing Inequality (DPI) to describe the information loss during processing [35]:
$I(Y;X)\geq I(Y;T)\geq I(Y;\hat{Y})$ (4)
In short, the generalization metric $I(Y;T)$ of hidden layer $T$ also describes the upper bound of $I(Y;\hat{Y})$, which is intrinsically a measure of robustness at layer $T$. ∎
@@
$||I^{*}(Y;T)-I(Y;T)||\leq\mathcal{O}(\frac{|\mathcal{T}||\mathcal{Y}|}{\sqrt{n}}).$ (6)
@@ 5 Experimental Results
To validate the Corollary 1, we evaluate the robustness in latent representations with respects to different depths by showcasing a small model VGG16 and a deeper model ResNet152. In both VGG16 and ResNet152, 5 feature extraction blocks which consist of several identical convolutional layers are used to extract features in different depths. We choose the first, the third and the last feature map denoted as Feature 0, Feature 2 and Feature 4, respectively. We choose PGD and Square Attack as the white- and black-box representatives respectively. To have a more comprehensive assessment, we explore the performance with multiple perturbation budget $\epsilon$.
As shown in Table 3, latent representations are always more robust than the input irrespective of the depth and perturbation strength, which is in line with our experiments in Section 5.1. Moreover, ASR decreases with the depth of feature maps for both VGG16 and ResNet152. Therefore, this experiment empirically reveal the trade-off between robustness and computational burden which is formally discussed in Corollary 1. Splitting DNNs in early layers can result in less computation in the mobile device but the latent space is more vulnerable to adversarial perturbations.
@@ 6 Related Work
Distributed Neural Networks. There is an increasing need for DNNs-based computation on mobile devices. Lightweight DNNs specifically tailored for mobile devices [4, 49, 50] fail to achieve comparable performance while edge computing approaches may incur in excessive latency [51]. As an intermediate option, [52] divide a DNN into two parts executed at the mobile device and edge, respectively. To reduce the networking load due to the large latent space of DNNs, literature introduces a “bottleneck” layer before the division point to reduce the dimension of latent representations [8, 11, 26]. Whereas, naive bottlenecks suffer from noticeable task-specific performance loss. As such, a key research problem in distributed DNN is to optimize the trade-off between the latent dimension and end-to-end performance. To achieve lower latent dimension while preserving accuracy, recent work utilizes more advanced training such as knowledge distillation [12] and multi-stage training [7]. Nevertheless, current work only focuses on the trade-off between the communication load and accuracy but ignore the robustness aspect.
In addition, different splitting points can affect the compression performance and computational burden in resource-constrained devices. For example, some studies [27, 53, 54] introduce bottlenecks in deeper layers and optimize a trainable rate-distortion metric. While such studies can achieve a remarkable compression performance, they neglect the limited computational capability of mobile devices which may result in larger delay for computation. In contrast, [28] introduce bottleneck in early layers with learnable prior for entropy coding to optimize the trade-off between compression and computation. [29] further investigate the effect of different splitting point in [28]. Our work reveals that the splitting point of DNN is highly related to the robustness to adversarial attacks in the distributed scenario.
@@
Empirical and Theoretical Study of DNN Robustness. [55] propose a set of criteria to study adversarial robustness with empirical results, while [56, 57, 58] evaluate the robustness of different defense approaches. However, in the absence of a theoretical framework, different studies may arrive at conflicting conclusions. For example, [38] argue that there is a trade-off between generalization and robustness while [59] state that generalization does not affect the robustness. [39] argue adversarial training sacrifice accuracy on standard datasets while [60] can achieve same accuracy on both adversarial and standard datasets. On the other hand, PAC learning has been used to analyze the adversarial robustness of DNNs [61, 62, 63, 64, 65, 66]. However, such work mainly attempts to find minimum dataset size for a desired robustness level [61, 63, 66, 62]. Conversely, our investigation aims to evaluate adversarial robustness in DNNs that are trained with certain datasets. In addition, these approaches are evaluated on simple DNNs [64, 65], while we evaluate our findings on state-of-the-art DNNs. [34] first propose to use IB [13] to analyze DNNs and [31, 41] analyze the generalization and compression capability of DNNs with experiments. However, due to the challenge of mutual information estimation [67], such studies only perform experiments on relatively small neural networks. Recent work applying IB for robust learning [68, 69, 70, 71] has shown significant enhancements. However, it does not involve a general analysis for robustness of DNNs.

### v5

@@ Abstract
Distributed deep neural networks (DNNs) have been shown to reduce the computational burden of mobile devices and decrease the end-to-end inference latency in edge computing scenariosApproved for Public Release: Distribution Unlimited: AFRL-2025-2602.. While distributed DNNs have been studied, the resilience of distributed DNNs to adversarial action remains an open problem. In this paper, we fill the existing research gap by rigorously analyzing the robustness of distributed DNNs against adversarial action. We cast this problem in the context of information theory and rigorously proved that (i) the compressed latent dimension improves the robustness but also affect task-oriented performance; and (ii) the deeper splitting point enhances the robustness but also increases the computational burden. These two trade-offs provide a novel perspective to design robust distributed DNN. To test our theoretical findings, we perform extensive experimental analysis by considering 6 different DNN architectures, 6 different approaches for distributed DNN and 10 different adversarial attacks using the ImageNet-1K dataset.
@@
Distributed deep neural network (DNN) , Adversarial Robustness , Split Computing
@@ 1 Introduction
DNNs have achieved significant success in various domains such as computer vision [1], natural language processing [2], and wireless communication [3], among many others. However, state-of-the-art DNNs are challenging to deploy on resource-limited mobile devices. While mobile-specific DNNs have been proposed [4], they usually come with a significant loss in accuracy. On the other hand, completely offloading the computation to edge or cloud computers is impractical in mobile scenarios due to the excessive communication overhead corresponding to the transfer of the DNN input from the mobile device to the edge/cloud [5]. A new paradigm called distributed computing – also referred to as split computing in prior art [6] – divides the computation of DNNs across multiple devices – according to the available processing power and networking bandwidth. The key advantage is that optimal load distribution can be achieved while meeting maximum end-to-end latency constraints and also preserving the DNN accuracy [7].
Without loss of generality, we assume that a DNN model is divided into a mobile DNN and a local DNN, respectively executed by the mobile device and an edge/cloud computer. Usually, the DNN architecture is modified by introducing a compression layer (“bottleneck”) at the end of the mobile DNN [8, 9], which is trained to learn a latent representation that reduces the amount of data being sent to the edge/cloud. The compressed representation is then used by the local DNN to produce the final prediction output (e.g., classification).
@@
Despite its significance and timeliness, assessing the robustness of distributed DNNs remains an unexplored problem. We remark that achieving a fundamental understanding of these attacks and evaluating their effectiveness in state-of-the-art DNNs is paramount to design robust distributed DNNs. To this end, we theoretically analyze the robustness of distributed DNNs using Information Bottleneck (IB) theory [10]. Our first key theoretical finding is that the latent robustness is highly depended on the depth of splitting point. With similar levels of information distortion, latent representations in deeper layers exhibit constantly better robustness compare to those in early layers. Our second key finding is that the trade-off between DNN performance and robustness is intrinsically related to the dimension of the latent space. While a smaller latent space may increase robustness by reducing the model variance, it will also introduce bias in the model thus affecting the generalization capability of the DNN model.
We extensively evaluate our theoretical findings by considering 10 adversarial algorithms, i.e., 4 white-box attacks [11, 12, 13, 14] and 6 black-box attacks [15, 16, 17, 18, 19, 20]. We apply these attacks to 6 different architectures [21, 22] designed with 6 distributed DNN approaches [23, 24, 9, 7, 25, 26]. The experimental results validate our theoretical findings on the examined DNNs and attack algorithms.
@@
$\bullet$ We perform extensive experiments with the ImageNet-1K [28] dataset, by considering 6 different DNN architectures, 6 different distributed DNN approaches under 10 different attacks to support our theoretical findings. The results show that the theoretical analysis applies to the experimental settings under consideration. We share our code for reproducibility at https://github.com/Restuccia-Group/AdvLatent, and we hope this work may open the door to a new field dedicated to studying the resilience of distributed DNNs.
@@
2 Overview of Distributed DNN
Distributed Neural Networks. Deploying AI applications in mobile devices is challenging as there is no enough computation resources to execute large DNNs. Lightweight DNNs [4, 29] have significant performance loss while offloading tasks to the edge device may incur in excessive latency [30]. Therefore, split computing, as an intermediate option, is proposed to accelerate large DNNs to resource-constrained devices [31]. The key idea is to split a large DNN into two parts – a relative small head model deployed on the resource-limited device and a large tail model deployed on the edge device which has excessive computation power.
“Partitioning optimization” and “bottleneck optimization” represent two major research directions in split computing. Partitioning optimization focuses on selecting the optimal splitting depth within DNNs to achieve minimal inference latency under given computation and bandwidth constraints [31, 23]. Splitting at early layers reduces computational overhead on the mobile device but increases communication burden, as latent representations in the initial layers of deep neural networks typically have larger dimensions. On the other hand, “bottleneck optimization” aims to minimize the size of latent representations by introducing a “bottleneck” layer before the split point to compress these representations [7, 32]. This approach involves a critical trade-off: smaller bottlenecks may cause information loss during compression, degrading end-to-end performance, while larger latent dimensions incur unnecessary communication overhead. For a comprehensive review, readers are referred to [33].
While the depth and dimension of the partitioning layer are key design factors in split computing, existing research has overlooked how these parameters affect the adversarial robustness of split computing systems. In this work, we investigate the relationship between adversarial robustness in the latent space and both the dimension and depth of partitioning layers.
Security and Privacy in Distributed DNNs. Despite its effectiveness, this new cooperative paradigm of distributed DNNs introduces significant security and privacy concerns that require investigation. We roughly categorize the security and privacy issues in distributed DNNs based on the target device. On mobile devices, a critical issue is data privacy. Literature has demonstrated that adversaries can extract private input information of the mobile device from the intermediate outputs [34]. In addition, research has shown that poisoning input data can compromise data compression mechanisms at the mobile side, creating excessive transmission overhead [35]. Conversely, we focus on a new security issue that targets the tail model deployed on the edge device.
@@ 3 Threat Model
where $\delta_{d}||_{p}\leq\sigma$ and $\sigma$ is the distance constraint under different $l_{p}$ norm. Additionally, for visual applications, $\delta_{d}$ should satisfy the condition $x+\delta_{d}\in[0,1]^{d}$ as there is an explicit upper and lower bound for red, green and blue (RGB) value in digital images.
@@
where $\delta_{t}||_{p}\leq\sigma$ is the distance constraint under $l_{p}$ norm. We remark that the latent representations are model-dependent and there is no explicit bound for their value other than their computer-level representation (e.g., float, integer, double).
@@ 4.1 Analysis through Information Bottleneck
The end-to-end robustness of the DNN model can be modeled with I(Y;Y^)I($\hat{Y}$), which measures the mutual information between the label $Y$ and the DNN inference $\hat{Y}$ [39]. We apply the Data Processing Inequality (DPI) to describe the information loss during processing [40]:
[…]
In short, the generalization metric $I(Y;T)$ of hidden layer $T$ also describes the upper bound of I(Y;Y^)I($\hat{Y}$), which is intrinsically a measure of robustness at layer $T$. ∎
@@
$I^{*}(Y;T)-I(Y;T)||\leq\mathcal{O}(\frac{|\mathcal{T}||\mathcal{Y}|}{\sqrt{n}}).$ (6)
@@ 6 Experimental Results
To validate Corollary 1, we evaluate the robustness of latent representations at different depths using two architectures: VGG16 as a shallow model and ResNet152 as a deeper model. Both architectures contain five feature extraction blocks, each comprising multiple convolutional layers that extract features at different depths. We select three splitting points for our analysis: the outputs of the first, third, and fifth blocks (denoted as Feature 0, Feature 2, and Feature 4, respectively). For a comprehensive evaluation, we employ PGD as a white-box attack and Square Attack as a black-box attack, testing multiple perturbation budgets $\epsilon$.
Table 3 demonstrates that latent representations consistently exhibit greater robustness than input-level perturbations, regardless of splitting depth or perturbation strength—corroborating our findings in Section 6.1. Notably, the ASR decreases monotonically with feature depth for both VGG16 and ResNet152, indicating that deeper splitting points yield more robust distributed DNNs.
To quantify the computational trade-off associated with this robustness gain, we deployed the distributed DNNs on a realistic edge computing setup: a Jetson Orin Nano serves as the mobile device executing the head model, while a NVIDIA A100 acts as the edge server executing the tail model. Table 4 presents the latency profiling results for batch processing of 32 inputs. Each row corresponds to a specific architecture and splitting configuration, with the “Head” and “Tail” columns reporting execution times on the mobile device and edge server, respectively.
The results reveal a clear trend: deeper splitting points increase the head execution latency, reflecting the increased computational burden on the mobile device. This empirical evidence demonstrates a fundamental trade-off in distributed DNN design. Splitting at later layers enhances robustness against latent-space adversarial perturbations but incurs higher computational costs on resource-constrained mobile devices. Conversely, early-layer splitting minimizes the mobile computational load but leaves the system more vulnerable to adversarial attacks. These findings empirically validate the theoretical framework presented in Corollary 1, highlighting the need for application-specific optimization when designing distributed inference systems.
@@
Table 4: Latency profiling of DNNs as a function of depth
Head (ms) Tail (ms) Comm. (ms) Overall (ms)
Feature 0 174.03 $\pm$ 17.11 13.42 $\pm$ 0.14 2211.61 $\pm$ 21.87 2399.06 $\pm$ 27.77
VGG16 Feature 2 386.77 $\pm$ 16.16 5.32 $\pm$ 0.06 63.50 $\pm$ 1.80 455.59 $\pm$ 16.26
Feature 4 525.59 $\pm$ 17.90 0.84 $\pm$ 0.01 1.92 $\pm$ 0.09 528.35 $\pm$ 17.90
Feature 0 24.22 $\pm$ 2.88 23.33 $\pm$ 0.09 67.66 $\pm$ 2.14 115.21 $\pm$ 3.59
ResNet152 Feature 2 217.57 $\pm$ 4.80 15.80 $\pm$ 0.07 458.26 $\pm$ 5.53 691.63 $\pm$ 7.32
Feature 4 611.26 $\pm$ 4.85 0.03 $\pm$ 0.00 0.12 $\pm$ 0.03 611.41 $\pm$ 4.85
6.5 Evaluation on Other Tasks
In this experiment, we provide a comprehensive assessment of our theoretical findings across different tasks. We consider two datasets: RAVDESS [54] which is commonly used in speech emotion recognition, and RadioML [55] for automatic modulation classification. For both tasks we use the same neural network architecture which is summarized in Table 5. For each conv layer, it follows by a batch normalization and ReLU activation.
Table 5: Summary of the DNN Classifier
[…]
Avgpool
[…]
For RAVDESS, we use the complete dataset and split it into training and testing sets with an 8:2 ratio. We employ an SGD optimizer with a learning rate of 0.01, weight decay of 5e-4, and momentum of 0.9. The model is trained for 200 epochs with a batch size of 32. Additionally, since RAVDESS is a small dataset, we apply label smoothing to prevent overfitting.
For RadioML, we use only data with SNR ranging from 10 dB to 20 dB, as signals at low SNR are challenging for classifiers to identify. We split the dataset into training and testing sets using the same configuration as RAVDESS. We employ the Adam optimizer with a learning rate of 0.001 and train for 50 epochs with a batch size of 1024. We do not apply label smoothing since RadioML is a large dataset.
After training, we split the model at the 4th conv layer and apply FGSM with $\epsilon=0.01$ to both input and latent space. Table 6 summarizes the classification accuracy on clean and adversarial data. The accuracy on latent adversarial samples consistently outperforms the performance on input adversarial samples across all datasets. This supports our theoretical analysis in Section 4.
Table 6: Input vs Latent Attack for Different Datasets
Dataset Clean Atk (input) Atk (latent)
RAVDESS [54] 82.25% 45.73% 65.87%
RadioML [55] 92.51% 58.34% 60.68%
@@ 7 Related Work
Adversarial Attacks. Adversarial attacks can be categorized as gradient-based, score-based, and decision-based. In gradient-based scenarios, attackers can obtain the input gradient through backpropagation and craft adversarial samples with gradient steps. FGSM [11] crafts adversarial samples in the $\mathit{l_{~\infty}}$ space based on the one-step input gradient sign. BIM [12] increases the effectiveness of FGSM by iteratively updating adversarial samples over multiple gradient steps. MIM [13] introduces momentum to iterative attacks which improves the transferability of adversarial samples. PGD [14] generalizes iterative attacks to $\mathit{l_{~p}}$ space with a random start. Score-based adversaries can only access the scores for every class given by the DNN. NES [15] applies evolutionary algorithm to estimate gradient within limit queries. N-Attack [16] designs a learnable Gaussian distribution to generate effective perturbations. Square Attack [17] adds a localized square-shaped perturbation at a random position to the input in each iteration. Decision-based attacks assume the adversary is only aware of the label having the highest score in the DNN output. EVO [18] minimizes perturbations with heuristic search. S-OPT [19] accelerates the convergence with estimated gradient sign, while Triangle Attack [20] minimizes perturbations in a low frequency space with the geometric property in the decision boundary.
@@
Empirical and Theoretical Study of DNN Robustness. [56] propose a set of criteria to study adversarial robustness with empirical results, and [57, 58] evaluate the robustness of different defense approaches. However, in the absence of a theoretical framework, different studies may arrive at conflicting conclusions. For example, [43] argue that there is a trade-off between generalization and robustness while [59] state that generalization does not affect the robustness. On the other hand, PAC learning has been used to theoretically analyze the adversarial robustness of DNNs [60, 61]. However, these approaches are evaluated on simple DNNs, while we evaluate our findings on state-of-the-art DNNs.
Information Bottleneck. [39] first propose to use IB [10] to analyze DNNs and [36, 46] analyze the generalization and compression capability of DNNs with experiments. However, due to the challenge of mutual information estimation [62], such studies only perform experiments on relatively small neural networks. Recent work applying IB for robust learning [63, 64] has shown significant enhancements in experimental results. However, it does not involve a rigorous theoretical analysis for robustness of DNNs. In contrast, we provide a theoretical framework to analyze the robustness of distributed DNNs in split computing as well as provide a comprehensive experimental assessment in large scale.
8 Limitation and Future Endeavor
Mutual Information Estimation. We employ IB as our theoretical framework. However, there has been extensive debate regarding whether IB constitutes an appropriate tool for explaining and interpreting the properties of DNNs [36, 46], as accurately quantifying mutual information in high-dimensional random variables presents significant challenges. In this work, we present a small-scale mutual information estimation experiment in Section 4.2 and evaluate the large scale experiment in Section 6 using end-to-end performance metrics rather than mutual information. Thus, this work can be more rigorous with advanced mutual information experiments in larger scale.
Generalization of Theoretical Findings. Our experiments focus exclusively on classification tasks. Investigating the framework’s applicability to other non-classification tasks would be valuable for future work. Furthermore, our experiments are performed on conventionally trained neural networks while evaluation on adversarially trained models [14] or dynamically robust models [65] remains unexplored.
Joint Optimization of Efficiency and Robustness. Note that the primary goal of this paper is to establish a fundamental understanding of robustness in the latent space of split computing. An important future direction is to develop an optimization framework that jointly optimizes task performance, end-to-end latency, and adversarial robustness in the latent space.
@@ 9 Concluding Remarks
Acknowledgment
This work has been supported in part by the National Science Foundation under grants CNS-2312875 and OAC-2530896; by the Air Force Office of Scientific Research under grant FA9550-23-1-0261; by the Office of Naval Research under grant N00014-23-1-2221.